# Artificial social influence via human-embodied AI agent interaction in immersive virtual reality (VR): Effects of similarity-matching during health conversations

### v1

@@ Abstract
Interactions with artificial intelligence (AI) based agents can positively influence human behavior and judgment. However, studies to date focus on text-based conversational agents (CA) with limited embodiment, restricting our understanding of how social influence principles, such as similarity, apply to AI agents (i.e., artificial social influence). We address this gap by leveraging the latest advances in AI (language models) and combining them with immersive virtual reality (VR). Specifically, we built VR-ECAs, or embodied conversational agents that can naturally converse with humans about health-related topics in a virtual environment. Then we manipulated human-agent similarity via gender matching and examined its effects on biobehavioral (i.e., gaze), social (e.g., agent likeability), and behavioral outcomes (i.e., healthy snack selection). We found that discussing health with opposite-gender agents enhanced gaze duration and the likelihood of healthy snack selection. In addition, female participants liked the VR-ECAs more than their male counterparts, regardless of the VR-ECAs’ gender. Finally, participants experienced greater presence while conversing with VR-embodied agents than chatting with text-only agents. Overall, our findings highlight embodiment as a crucial factor of AI’s influence on human behavior, and our paradigm enables new experimental research at the intersection of social influence, human-AI communication, and immersive virtual reality (VR).
@@ 1 Introduction
This study introduces an innovative research paradigm that involves real-time conversations with VR-ECAs about health. Theoretically, we examine the effect of human-agent similarity on participants’ gaze towards the agent, evaluations (e.g., agent likeability), and behavior (i.e., snack selection). In this paper, we first introduce each aspect of our research paradigm. Then we build our research questions and hypotheses based on existing work. Next, we outline the methodology and present the results of our experimental manipulations. Finally, we discuss how the findings advance our understanding of artificial influence and computer-mediated communication more broadly.
@@ 1.2 Virtual Reality (VR): Embodiment, Immersive Real-time Interactions, and Biobehavioral Measurement
The concept of immersiveness refers to the capability of technology to absorb users’ perceptual system, block out real-world sensory input, and stimulate it authentically with artificially generated simulacra (Bente et al., 2023; Biocca and Delaney, 1995). VR’s immersive nature thus offers the unique opportunity for humans to encounter artificial agents in the same virtual space and engage in verbal back-and-forth dialogue with them. With biobehavioral measures like eye-tracking, which are increasingly integrated into head-mounted displays, we can now gather more precise information about the mechanisms of interpersonal interaction within virtual environments (Andrei et al., 2023; Syrjämäki et al., 2020). Indeed, scholars have already begun to examine how factors like human-to-agent similarity influence the interaction (Shih et al., 2023), and we expect a rapid expansion of VR-based research with artificial agents.
@@ 1.3 Similarity: Essential Ingredient of Artificial Influence
It is well-known that similarity among interacting individuals - both observable and perceived - has a powerful effect on human behavior. For instance, human relationships and connections are often formed based on shared attributes, including ethnic and cultural backgrounds, interests, and attitudes (i.e., homophily; McPherson et al. (2001)). Rogers and Bhowmik (1970) additionally proposed that - when given a choice - people prefer to engage with and become more receptive to those who resemble them. Likewise, Cialdini and Goldstein (2004) suggested that during interactions with strangers, similarity acts as a heuristic for potential future friendship or acquaintance, enhancing the likelihood of compliance. Social identity theory (SIT) further underscores the link between similarity and favorable outcomes. SIT posits that people are driven to categorize themselves and others into distinct groups (based on factors such as gender, nationality, political or ideological affiliations, etc.), and think and act according to the group they identify with (Ashforth and Mael, 1989; Hogg et al., 1995; Tajfel and Turner, 2004). Thus, messages from those categorized as similar (i.e., in-group members) tend to exhibit greater influence than messages from dissimilar individuals (i.e., out-group members; Wood (2000)). The influence of similarity has also been demonstrated in human-agent interaction literature: Studies found that similarity enhanced social outcomes (e.g., perception of chatbot friendliness; Jin and Eastin (2023); interpersonal closeness; Liao and He (2020); enjoyability and sociability; Qiu and Benbasat (2010)) and favorable outcomes (e.g., favorable attitudes; Jin and Eastin (2023)).
Building on these findings, this study examined the effect of human-agent similarity (via gender matching) on agent likeability and healthy food choice after a natural dialogue about health in an immersive VR environment (see Figure 1). Specifically, we created two types of AI health coaches. The first type was a VR-ECA, which integrated OpenAI GPT4 and text-to-speech and speech-to-text AI models with avatars that displayed basic lip sync and gaze behavior in a virtual reality platform. The second type was a text-based CA with no embodiment cues, created via the OpenAI platform. In a between-subjects design, participants were randomly assigned to interact with either a gender-matched VR-ECA, a gender-unmatched VR-ECA, or with the text-based CA. Our study examined two important questions: 1) the effect of immersive VR and embodiment on perceived presence and 2) the effect of gender matching on social (e.g., agents’ perceived likeability), behavioral (i.e., choosing a healthy snack), and biobehavioral outcomes (i.e., gaze duration).
@@ 3.1 Participants
We recruited 60 participants via a university pool for research credit and word of mouth to broaden the sample demographics. The institutional review board approved the study. All participants were included in the final sample (21 gender-matched VR-ECA, 19 gender-unmatched VR-ECA, and 20 text-based CA; mage = 23.33; sdage = 8.25), with 63% identifying as female and 60% identifying as White or European American. Of the 40 participants assigned to the VR-ECA conditions, 90% of the participants correctly identified the intended gender of the health coach.
@@ 3.4.1 Perceived Presence
We adopted the perceived presence scale from Bente et al. (2023). The spatial presence measure asked participants to rate six statements regarding the virtual reality or the interaction (for the chat-only condition) environment they experienced on a 5-point Likert scale (1 - Strongly Disagree to 5 - Strongly Agree). The measure showed a good reliability score of .89 across the two tasks (MConsultation = 3.16, SDConsultation = .93). The co-presence scale comprised six items in the same 5-Likert scale related to the sense of being in the same place as the AI health coach, and the items exhibited good reliability (Cronbach’s alphaConsultation = .85; MConsultation = 3.19, SDConsultation = .88). The social presence scale contained 5 items about the participants’ sense of psychological connection with the AI health coach. These 5-point Likert scale items had adequate reliability on average (Cronbach’s alphaConsultation = .71; MConsultation = 3.75, SDConsultation = .71).
@@ 4.2 RQ2, H1-H2: Effect of Gender Matching on Gaze, Likeability, and Snack Selection
Next, we present the results from a series of regression models applied to the VR-ECA conditions (see Figure 4 and Table 2). We first found a significant interaction effect between the users’ gender identification and the agents’ gender (F(1, 33) = 6.85, p = .013) on perceived appearance similarity. The pairwise comparisons showed that male participants felt more similar to male VR-ECAs than female VR-ECAs (difference = -1.61, SE = .72, p = .033; see Table 3). In addition, female participants felt more similar to female VR-ECAs compared to their male participant counterparts (difference = 2.14, SE = .70, p = .0042). This suggested that our gender matching manipulation was successful and induced intended effects on perceived similarity to the agent’s appearance. Gender matching did not affect perceived similarity in attitude (F(1, 33) = .077, p = .78).
@@
Intercept 3.78 .46 t = 8.28 $<$.001
VR-ECA Gender: Female -1.61 .72 t = -2.23 .033
User Gender: Female -.29 .62 t = -.48 .64
VR-ECA Gender: Female x User Gender: Female 2.43 .93 t = 2.62 .013
@@
User Gender: Female .85 .60 t = 1.41 .17
VR-ECA Gender: Female x User Gender: Female -.25 .91 t = -.28 .78
@@
User Gender: Female 1.43 .47 t = 3.03 .0047
VR-ECA Gender: Female x User Gender: Female -.35 .71 t = -.49 .63
@@
User Gender: Female 1.73 1.03 z = 1.68 .094
VR-ECA Gender: Female x User Gender: Female -3.52 1.62 z = -2.17 .030
@@
Intercept -.05 .23 z = -.22 .82
VR-ECA Gender: Female .97 .37 z = 2.59 .0096
User Gender: Female 1.07 .32 z = 3.35 $<$.001
VR-ECA Gender: Female x User Gender: Female -1.42 .48 z = -2.95 .0032
@@
Male VR-ECA + Male User vs. Male VR-ECA + Female User .29 .62 t = .48 .64
Female VR-ECA + Female User vs. Female VR-ECA + Male User 2.14 .70 t = 3.07 .0042
Male VR-ECA + Male User vs. Female VR-ECA + Male User 1.61 .72 t = 2.23 .033
Female VR-ECA + Female User vs. Male VR-ECA + Female User .82 .58 t = 1.40 .17
@@
Male VR-ECA + Male User vs. Male VR-ECA + Female User -1.43 .47 t = -3.03 .0047
Female VR-ECA + Female User vs. Female VR-ECA + Male User 1.08 .53 t = 2.03 .05
@@
Female VR-ECA + Female User vs. Male VR-ECA + Female User -.61 .45 t = -1.36 .18
@@
Male VR-ECA + Male User vs. Male VR-ECA + Female User -1.73 1.03 t = -1.68 .094
Female VR-ECA + Female User vs. Female VR-ECA + Male User -1.79 1.25 t = -1.43 .15
@@
Female VR-ECA + Female User vs. Male VR-ECA + Female User -1.69 .99 t = -1.71 .089
@@
Male VR-ECA + Male User vs. Male VR-ECA + Female User -.25 .071 z = -3.46 $<$.001
Female VR-ECA + Female User vs. Female VR-ECA + Male User -.079 .077 z = -1.02 .31
@@
Female VR-ECA + Female User vs. Male VR-ECA + Female User -.098 .065 z = -1.51 .13
@@
For gaze ratio, there was a significant interaction between VR-ECA and participant gender ($\chi^{2}$ = 8.69, p = .0032). Pairwise comparisons showed that female participants looked at male VR-ECAs longer than their male counterparts (difference = -.25, SE = .071, p < .001), and male participants looked at female VR-ECAs longer than the male VR-ECAs (difference = -.23, SE = .083, p = .0061).
Snack selection illustrated a similar occurrence as gaze ratio: there was a significant interaction effect between the user and VR-ECA’s gender ($\chi^{2}$ = 5.50, p = .019), but in the opposite direction than we expected. Pairwise comparisons showed that people were (directionally) more likely to select the healthy snack after interacting with the VR-ECA of the opposite gender. Thus, H1 was not supported.
Contrary to our predictions, gender matching did not increase the likeability of the VR-ECAs though (F(1, 33) = .24, p = .63). However, we found a significant main effect of participant gender; female participants liked their agents more than their male counterparts, regardless of the VR-ECAs’ gender (F(1, 33) = 9.17, p = .0047). Therefore, H2 was not supported.
@@ 5 Discussion
Next, this study provided interesting insights into the effect of human-agent similarity. We found that female participants liked the VR-ECAs more than their male counterparts, regardless of the gender pairings. Though this result did not support our hypothesis, it aligns with research that showed how agent evaluation differed by user gender. For instance, female participants in Liu and Yao (2023) were evaluated as being nicer and more pleasant to interact with while interacting with Amazon’s Alexa VoiceBot than their male counterparts. Similarly, female participants in Krämer et al. (2010) evaluated an embodied agent more favorably than male participants. These results suggest that users’ gender identity could influence people’s responses to technology and media (Brunel and Nelson, 2003; Cai et al., 2017; Darley and Smith, 1995; Deaux, 1984).
Regarding the effect of similarity on snack selection and gaze, we found that opposite-gender pairings actually enhanced outcomes. For gaze ratio, participants looked more at the opposite-gender agent during the conversation compared to same-gender agents. In addition, participants were more likely to select a healthy snack after the conversation if the VR-ECA was the opposite gender. While few studies have examined the influence of dissimilarity, Zanbaka et al. (2006) showed that interacting with virtual speakers of the opposite gender improved attitudes toward the topics of persuasive messages. Participants in Carli (1990) also reported that the opposite-gender (human) speakers had more influence. These results suggest that factors other than similarity influenced outcomes. While more research is needed to fully unpack and replicate these results, we offer a few potential explanations about why gender matching did not elicit expected outcomes. One explanation relates to social scripts, or social and biological factors related to attraction and mate selection. Although participants were certainly aware of the fact that they were interacting with an AI agent, the visual embodiment created a very immediate impression that may have steered automatic social behaviors. For instance, both male as well as female agents were generally considered attractive, having a symmetric face, clear skin, and a well-proportioned body in line with gender-typical physical attractiveness cues (e.g. man: strong jaw, muscular body; woman: full lips and high waist-to-hip ratio). Thus, given that gaze behavior is largely implicit and beyond conscious control, participants might have been more interested in, and attentive to the other gender. This effect would run counter to our similarity manipulation. Also, participants would become more health conscious to appear more attractive to the opposite gender. This biological response would run counter to the intended similarity manipulation.

### v2

@@ Abstract
Interactions with artificial intelligence (AI) based agents can positively influence human behavior and judgment. However, studies to date focus on text-based conversational agents (CA) with limited embodiment, restricting our understanding of how social influence principles, such as similarity, apply to AI agents (i.e., artificial social influence). We address this gap by leveraging the latest advances in AI (language models) and combining them with immersive virtual reality (VR). Specifically, we built VR-ECAs, or embodied conversational agents that can naturally converse with humans about health-related topics in a virtual environment. Then we manipulated interpersonal similarity via gender matching and examined its effects on biobehavioral (i.e., gaze), social (e.g., agent likeability), and behavioral outcomes (i.e., healthy snack selection). We found an interesting interaction effect between agent and participant gender on biobehavioral outcomes: discussing health with opposite-gender agents tended to enhance gaze duration, with the effect stronger for male participants compared to their female counterparts. A similar directional pattern was observed for healthy snack selection, though it was not statistically significant. In addition, female participants liked the VR-ECAs more than their male counterparts, regardless of the VR-ECAs’ gender. Finally, participants experienced greater presence while conversing with VR-embodied agents than chatting with text-only agents. Overall, our findings highlight embodiment as a crucial factor of AI’s influence on human behavior, and our paradigm enables new experimental research at the intersection of social influence, human-AI communication, and immersive virtual reality (VR).
@@ 1 Introduction
This study introduces an innovative research paradigm that involves real-time conversations with VR-ECAs about health. Theoretically, we examine the effect of interpersonal similarity on participants’ gaze towards the agent, evaluations (e.g., agent likeability), and behavior (i.e., snack selection). In this paper, we first introduce each aspect of our research paradigm. Then we build our research questions and hypotheses based on existing work. Next, we outline the methodology and present the results of our experimental manipulations. Finally, we discuss how the findings advance our understanding of artificial influence and computer-mediated communication more broadly.
@@ 1.2 Virtual Reality (VR): Embodiment, Immersive Real-time Interactions, and Biobehavioral Measurement
The concept of immersiveness refers to the capability of technology to absorb users’ perceptual system, block out real-world sensory input, and stimulate it authentically with artificially generated simulacra (Bente et al., 2023; Biocca and Delaney, 1995). VR’s immersive nature thus offers the unique opportunity for humans to encounter artificial agents in the same virtual space and engage in verbal back-and-forth dialogue with them. With biobehavioral measures like eye-tracking, which are increasingly integrated into head-mounted displays, we can now gather more precise information about the mechanisms of interpersonal interaction within virtual environments (Andrei et al., 2023; Syrjämäki et al., 2020). Indeed, scholars have already begun to examine how factors like interpersonal similarity influence the interaction (Shih et al., 2023), and we expect a rapid expansion of VR-based research with artificial agents.
@@ 1.3 Similarity: Essential Ingredient of Artificial Influence
It is well-known that similarity among interacting individuals - both observable and perceived - has a powerful effect on human behavior. For instance, human relationships and connections are often formed based on shared attributes, including ethnic and cultural backgrounds, interests, and attitudes (i.e., homophily; (McPherson et al., 2001)). Rogers and Bhowmik (1970) additionally proposed that - when given a choice - people prefer to engage with and become more receptive to those who resemble them. Likewise, Cialdini and Goldstein (2004) suggested that during interactions with strangers, similarity acts as a heuristic for potential future friendship or acquaintance, enhancing the likelihood of compliance. Social identity theory (SIT) further underscores the link between similarity and favorable outcomes. SIT posits that people are driven to categorize themselves and others into distinct groups (based on factors such as gender, nationality, political or ideological affiliations, etc.), and think and act according to the group they identify with (Ashforth and Mael, 1989; Hogg et al., 1995; Tajfel and Turner, 2004). Thus, messages from those categorized as similar (i.e., in-group members) tend to exhibit greater influence than messages from dissimilar individuals (i.e., out-group members; Wood (2000)). The influence of similarity has also been demonstrated in human-agent interaction literature: Studies found that similarity enhanced social outcomes (e.g., perception of chatbot friendliness; Jin and Eastin (2023); interpersonal closeness; Liao and He (2020); enjoyability and sociability; Qiu and Benbasat (2010)) and favorable outcomes (e.g., favorable attitudes; Jin and Eastin (2023)).
Building on these findings, this study examined the effect of interpersonal similarity (via gender matching) on agent likeability and healthy food choice after a natural dialogue about health in an immersive VR environment (see Figure 1). Specifically, we created two types of AI health coaches. The first type was a VR-ECA, which integrated OpenAI GPT4 and text-to-speech and speech-to-text AI models with avatars that displayed basic lip sync and gaze behavior in a virtual reality platform. The second type was a text-based CA with no embodiment cues, created via the OpenAI platform. In a between-subjects design, participants were randomly assigned to interact with either a gender-matched VR-ECA, a gender-unmatched VR-ECA, or with the text-based CA. Our study examined two important questions: 1) the effect of immersive VR and embodiment on perceived presence and 2) the effect of gender matching on social (e.g., agents’ perceived likeability), behavioral (i.e., choosing a healthy snack), and biobehavioral outcomes (i.e., gaze duration).
@@ 3.1 Participants
We recruited 60 participants via a university pool for research credit and word of mouth to broaden the sample demographics. The institutional review board approved the study. All participants were included in the final sample (19 gender-matched VR-ECA, 21 gender-unmatched VR-ECA, and 20 text-based CA; mage = 23.23; sdage = 8.25), with 63% identifying as female and 60% identifying as White or European American. Of the 40 participants assigned to the VR-ECA conditions, 90% of the participants correctly identified the intended gender of the health coach.
@@ 3.4.1 Perceived Presence
We adopted the perceived presence scale from Bente et al. (2023). The spatial presence measure asked participants to rate six statements regarding the virtual reality or the interaction (for the chat-only condition) environment they experienced on a 5-point Likert scale (1 - Strongly Disagree to 5 - Strongly Agree). The measure showed a good reliability score of .89 across the two tasks (MConsultation = 3.16, SDConsultation = .93). The co-presence scale comprised six items in the same 5-Likert scale related to the sense of being in the same place as the AI health coach, and the items exhibited good reliability (Cronbach’s alphaConsultation = .85; MConsultation = 3.19, SDConsultation = .88). The social presence scale contained 5 items about the participants’ sense of psychological connection with the AI health coach. These 5-point Likert scale items had adequate reliability on average (Cronbach’s alphaConsultation = .71; MConsultation = 3.05, SDConsultation = .65).
@@ 4.2 RQ2, H1-H2: Effect of Gender Matching on Gaze, Likeability, and Snack Selection
Next, we present the results from a series of regression models applied to the VR-ECA conditions (see Figure 4 and Table 2). We first found a significant interaction effect between the users’ gender identification and the agents’ gender (F(1, 33) = 5.62, p = .024) on perceived appearance similarity. The pairwise comparisons showed that male participants felt more similar to male VR-ECAs than female VR-ECAs (difference = -1.61, SE = .73, p = .035; see Table 3). In addition, female participants felt more similar to female VR-ECAs compared to their male participant counterparts (difference = 2.07, SE = .72, p = .0068). This suggested that our gender matching manipulation was generally successful. Gender matching did not affect perceived similarity in attitude (F(1, 33) = .17, p = .69).
@@
Intercept 3.78 .46 t = 8.17 $<$.001
VR-ECA Gender: Female -1.61 .73 t = -2.20 .035
User Gender: Female -.17 .61 t = -.27 79
VR-ECA Gender: Female x User Gender: Female 2.23 .94 t = 2.37 .024
@@
User Gender: Female .90 .59 t = 1.53 .14
VR-ECA Gender: Female x User Gender: Female -.37 .91 t = -.41 .69
@@
User Gender: Female 1.41 .46 t = 3.05 .0045
VR-ECA Gender: Female x User Gender: Female -.36 .71 t = -.51 .61
@@
User Gender: Female 1.32 .95 z = 1.40 .16
VR-ECA Gender: Female x User Gender: Female -2.93 1.58 z = -1.86 .064
@@
Intercept -.05 .22 z = -.22 .82
VR-ECA Gender: Female .97 .37 z = 2.59 .0095
User Gender: Female 1.05 .31 z = 3.37 $<$.001
VR-ECA Gender: Female x User Gender: Female -1.43 .48 z = -2.97 .0030
@@
Male VR-ECA + Male User vs. Male VR-ECA + Female User .17 .61 t = .27 .79
Female VR-ECA + Female User vs. Female VR-ECA + Male User 2.07 .72 t = 2.89 .0068
Male VR-ECA + Male User vs. Female VR-ECA + Male User 1.61 .73 t = 2.20 .035
Female VR-ECA + Female User vs. Male VR-ECA + Female User .62 .59 t = 1.05 .30
@@
Male VR-ECA + Male User vs. Male VR-ECA + Female User -1.41 .46 t = -3.05 .0045
Female VR-ECA + Female User vs. Female VR-ECA + Male User 1.04 .54 t = 1.93 .062
@@
Female VR-ECA + Female User vs. Male VR-ECA + Female User -.62 .45 t = -1.39 .17
@@
Male VR-ECA + Male User vs. Male VR-ECA + Female User -1.32 .95 t = -1.40 .16
Female VR-ECA + Female User vs. Female VR-ECA + Male User -1.61 1.27 t = -1.27 .20
@@
Female VR-ECA + Female User vs. Male VR-ECA + Female User -1.10 .92 t = -1.20 .23
@@
Male VR-ECA + Male User vs. Male VR-ECA + Female User -.24 .070 z = -3.46 $<$.001
Female VR-ECA + Female User vs. Female VR-ECA + Male User -.084 .079 z = -1.07 .29
@@
Female VR-ECA + Female User vs. Male VR-ECA + Female User -.10 .066 z = -1.53 .13
@@
For gaze ratio, there was a significant interaction between VR-ECA and participant gender ($\chi^{2}$ = 8.80, p = .0030). Pairwise comparisons showed that male participants looked at female VR-ECAs longer than the male VR-ECAs (difference = -.23, SE = .083, p = .0061). Female participants looked at male VR-ECAs longer than their male counterparts (difference = -.24, SE = .070, p < .001). Snack selection illustrated a similar occurrence as gaze ratio: there was a significant interaction effect between the user and VR-ECA’s gender ($\chi^{2}$ = 3.90, p = .048). Figure 4 suggests that people who interacted with the VR-ECA of the opposite gender slightly trended toward selecting a healthier snack, but the result was not statistically significant. Thus, H1 was not supported.
Interestingly, gender matching did not effect the likeability of the VR-ECAs (F(1, 33) = .26, p = .61). However, we found a significant main effect of participant gender; female participants liked their agents more than their male counterparts, regardless of the VR-ECAs’ gender (F(1, 33) = 9.29, p = .0045). Therefore, H2 was not supported.
@@ 5 Discussion
Next, this study provided interesting insights into the effect of interpersonal similarity. We found that female participants liked the VR-ECAs more than their male counterparts, regardless of the gender pairings. Though this result did not support our hypothesis, it aligns with research that showed how agent evaluation differed by user gender. For instance, female participants in Liu and Yao (2023) were evaluated as being nicer and more pleasant to interact with while interacting with Amazon’s Alexa VoiceBot than their male counterparts. Similarly, female participants in Krämer et al. (2010) evaluated an embodied agent more favorably than male participants. These results suggest that users’ gender identity could influence people’s responses to technology and media (Brunel and Nelson, 2003; Cai et al., 2017; Darley and Smith, 1995; Deaux, 1984).
Regarding the effect of similarity on gaze and snack selection, we found that opposite-gender pairings actually showed potential to enhance outcomes. For gaze ratio, male participants looked more at the female VR-ECA than the male VR-ECAs during the conversation and female participants looked more at the male VR-ECAs than their male counterparts. In addition, participants demonstrated slight trend of selecting a healthy snack after the conversation if the VR-ECA was the opposite gender (though the results were not statistically significant). While few studies have examined the influence of dissimilarity, Zanbaka et al. (2006) showed that interacting with virtual speakers of the opposite gender improved attitudes toward the topics of persuasive messages. Participants in Carli (1990) also reported that the opposite-gender (human) speakers had more influence. These results suggest that factors other than similarity influenced outcomes. While more research is needed to fully unpack and replicate these results, we offer a few potential explanations about why gender matching did not elicit expected outcomes. One explanation relates to social scripts, or social and biological factors related to attraction and mate selection. Although participants were certainly aware of the fact that they were interacting with an AI agent, the visual embodiment created a very immediate impression that may have steered automatic social behaviors. For instance, both male as well as female agents were generally considered attractive, having a symmetric face, clear skin, and a well-proportioned body in line with gender-typical physical attractiveness cues (e.g. man: strong jaw, muscular body; woman: full lips and high waist-to-hip ratio). Thus, given that gaze behavior is largely implicit and beyond conscious control, participants might have been more interested in, and attentive to the other gender. This effect would run counter to our similarity manipulation. Also, participants would become more health conscious to appear more attractive to the opposite gender. This biological response would run counter to the intended similarity manipulation.